OpenStudy (anonymous):

Help! Medal and Fan, if you can help! solve: f(t)=20+150*ln(8t+1) and I have to find 10 minutes, which will then give me the equation f(10)=20+150*ln(8(10)+1) Can someone help solve it?

OpenStudy (anonymous):

@YanaSidlinskiy This is the one!

OpenStudy (yanasidlinskiy):

Aww...What a problem..

OpenStudy (anonymous):

Yeah! I got f(10)=170*ln(81) but then i'm stuck and i don't even know if it's right..

OpenStudy (anonymous):

or at least it should be 20+150*ln(81)... but still stuck on what i should do

OpenStudy (anonymous):

ln= ? ? What do you mean?

OpenStudy (yanasidlinskiy):

f(10)=170*ln(81)

OpenStudy (anonymous):

That's how far I got, but i need to get the temp.

OpenStudy (yanasidlinskiy):

@phi Can you help?

OpenStudy (anonymous):

@AccessDenied, can you help with this?

OpenStudy (anonymous):

Or someone -.- @mathmale ?

OpenStudy (anonymous):

@whpalmer4 ?

OpenStudy (whpalmer4):

\[ f(t)=20+150*\ln(8t+1)\] \[f(10) = 20+150*\ln(8(10)+1)\]\[f(10) = 20+150*\ln(81)\]

OpenStudy (anonymous):

@whpalmer4 yes, i got to that part but can't i solve to get one number? I need to find the temp (f(t))

OpenStudy (whpalmer4):

So don't you have a calculator?

OpenStudy (anonymous):

No, actually...

OpenStudy (whpalmer4):

Or select values of \(\ln x\) committed to memory? :-) \[\ln 3 \approx 1.09861\]so \[\ln 81 = \ln 3^4 = 4\ln 3 \approx 4*1.09861 \approx 4.394\]

OpenStudy (whpalmer4):

http://wolframalpha.com is a fine calculator, as is http://web2.0calc.com/

OpenStudy (anonymous):

I've tried wolframalpha, that isn't working... But i'll try the other one. Thanks

OpenStudy (whpalmer4):

oh sure it is.... http://www.wolframalpha.com/input/?i=ln+81

OpenStudy (anonymous):

Ah... Okay. Thanks for answering! :)

OpenStudy (whpalmer4):

actually, it's the other one that isn't working now :-)

OpenStudy (whpalmer4):

if you just go to google.com and type your expressions into the search box, that usually works, too. https://www.google.com/#q=ln+81

OpenStudy (anonymous):

I haven't tried the other one yet... I put it in wrong in wolfram, that's why i couldn't get any answer.

OpenStudy (whpalmer4):

wolfram would probably interpret your problem almost directly: http://www.wolframalpha.com/input/?i=f(t)%3D20%2B150*ln(8t%2B1)+with+t%3D10

OpenStudy (anonymous):

This is awesome help! I'm looking through it now :D Thanks !!

OpenStudy (whpalmer4):

it's a great tool, but resist the temptation to have it do your homework!

OpenStudy (anonymous):

Lol, yup. Got you.. That's why I come here to get help so i can work with and understand it

OpenStudy (anonymous):

Find the interquartile range of the data. A. 6 B. 8 C. 9 D. 20

OpenStudy (whpalmer4):

@natedog12 do not post your questions in someone else's question!

OpenStudy (anonymous):

dude that was 30 minutes ago

OpenStudy (whpalmer4):

Doesn't matter, just don't do it.